Zinc metal reacts with hydrochloric acid to produce hydrogen gas and zinc chloride. Given a 0.215 g piece of zinc reacting with 57ml of 0.10m hydrochloric acid, determine the maximum volume of hydrogen gas that can be formed at 20 degrees Celsius and 760mmhg. (This is a limiting reactant problem)
1
Expert's answer
2019-04-11T06:00:48-0400

The reaction between zink metal and hydrochloric acid is as followong:

Zn + 2KCl = ZnCl2 + H2

The number of moles of zink:

n(Zn) = m(Zn) / Mr(Zn) = 0.215g / 65 g/mol = 0.0033 mol

The number of moles of hydrogen chloride:

n(HCl) = cM × V = 0.1 mol/L × 0.057 L = 0.0057 mol.

From 0.0033 mol of Zn, the same amounts of H2 can be produced.

However, from 0.0057 mol of HCl, only 0.0057 / 2 = 0.00285 mol of H2 can be produced. As a result, HCl is a limiting reactant and 0.00285 mol of H2 are produced in the reaction.

To find the volume of H2, the ideal gas equation can be used. From the ideal gas equation:

V = nRT/P,

where V - volume, n - number of moles, R - gas constant, T - temperature, P - pressure.

Here, n = 0.00285 mol, R = 0.0821 L·atm·mol-1·K-1, T = 20°C = 293.15 K, P = 760 mmHg = 1 atm.

As a result,

V(H2) = 0.00285 mol × 0.0821 L·atm·mol-1·K-1 × 293.15 K / 1 atm = 0.07 L = 70 mL.

Answer: 0.07L = 70 mL of H2.